OpenStudy (zion):

f(x)=X-4/x, and f(2b)=2f(b), then b is? solutions..?

OpenStudy (zion):

manoranjan! help

OpenStudy (akshay_budhkar):

see solve the equation 2x-4/2x=2[x-4/x]

OpenStudy (anonymous):

inconsistent, no?

OpenStudy (zion):

the answer in d reviewer is 6,.. but what is the solution to get 6?

OpenStudy (anonymous):

If b is 6 then f(6) = 5 1/3 *2 = 10 2/3 but f(12) is 11 2/3 so b is not 6.

OpenStudy (phi):

Zion, is this the right equation?? \[f(x)=X-\frac{4}{x}\] does capital X have a known value?? As written, there is no solution, so maybe there's a typo ?

OpenStudy (phi):

if on the other hand, the equation is \[f(x)= \frac{x-4}{x}\] we can solve it.

OpenStudy (zion):

thats the equation!

OpenStudy (zion):

how did u do that,.. coz i dont know how to type that equation here

OpenStudy (phi):

In the equation editor, you can type "\"frac (without the quotes!){top}{bottom}

OpenStudy (zion):

so whats the solution to my question?

OpenStudy (phi):

But back to the problem. If we use \[f(x)=\frac{x-4}{x}\] substitute 2b for x. Try it!

OpenStudy (zion):

i still dont know how

OpenStudy (phi):

to do what?

OpenStudy (zion):

what im going to do.

OpenStudy (phi):

OK, here's the big picture...

OpenStudy (anonymous):

Still not going to work for 6, is it?

OpenStudy (phi):

you have a function f(x) = (x-4)/x

OpenStudy (zion):

then...?

OpenStudy (phi):

This means wherever you see an x, you can replace it.. I mean if you have f(b), you replace the x with a b... or if you have f(2b), replace the x with 2b.

OpenStudy (phi):

So the very first thing to do is f(b) and then f(2b). Give it a shot.

OpenStudy (phi):

Here's f(b) f(b)= (b-4)/b do the same for f(2b). Then on to the next step

OpenStudy (phi):

Did I lose you?

OpenStudy (anonymous):

f(x)=(x-4)/x f(2b)=(2b-4)/2b f(2b)= 2 f(b) f(2b)=f(2b] (2b-4)/2b = 2(b-4)/b cross multiply b(2b-4) = 2b(2b-8) cancel b on both sides 2b-4 = 2(2b-8) 2b-4 = 4b-16 group similar terms 4b-2b = 16-4 2b = 12 b = 6

OpenStudy (phi):

Yes It looks good. No, it looks very good! But you aren't zion.